# John Mendonça1,2, Isabel Trancoso1,2    Alon Lavie3,4

### v3

@@ Abstract
Evaluating the quality of open-domain chatbots has become increasingly reliant on LLMs acting as automatic judges. However, existing meta-evaluation benchmarks are static, outdated, and lacking in multilingual coverage, limiting their ability to fully capture subtle weaknesses in evaluation. We introduce  [Uncaptioned image] MEDAL, an automated multi-agent framework for curating more representative and diverse open-domain dialogue evaluation benchmarks [^1] Code and data: github.com/johndmendonca/medal.. Our approach leverages several state-of-the-art LLMs to generate user-chatbot multilingual dialogues, conditioned on varied seed contexts. Then, a strong LLM (GPT-4.1) is used for a multidimensional analysis of the performance of the chatbots, uncovering noticeable cross-lingual performance differences. Guided by this large-scale evaluation, we curate a new meta-evaluation multilingual benchmark and human-annotate samples with nuanced quality judgments. This benchmark is then used to assess the ability of several reasoning and non-reasoning LLMs to act as evaluators of open-domain dialogues. Using MEDAL, we uncover that state-of-the-art judges fail to reliably detect nuanced issues such as lack of empathy, commonsense, or relevance.
@@
John Mendonça1,2, Isabel Trancoso1,2and Alon Lavie3,4 1 INESC-ID, Lisbon 2 Instituto Superior Técnico, University of Lisbon 3 Carnegie Mellon University, Pittsburgh 4 Phrase, Pittsburgh {john.mendonca, isabel.trancoso}@inesc-id.pt, alavie@cs.cmu.edu
@@ 1 Introduction
Figure 1: Overview of  Refer to caption MEDAL, a 3-step framework for curating multilingual meta-evaluation benchmarks.
Dialogue generation (§3): Generation of multi-turn, multilingual synthetic dialogues seeded from diverse contexts, between an LLM acting as a user (with utterances validated by a LLM-judge) and the target chatbot; ➋ Automated labelling (§4): Multidimensional evaluation of all dialogues by a strong LLM (GPT-4.1); ➌ Benchmark Curation (§5): Curation of meta-evaluation benchmarks using human annotations used to assess LLM judges.
@@
    Dialogue generation (§3): synthetically generate a large collection of dialogues in multiple languages, using several state-of-the-art LLMs acting as agents in the dialogue-generation process.
@@
    Benchmark curation (§5): guided by these automated labels, we curate and human-annotate a balanced set of dialogues to form a meta-evaluation benchmark that directly tests LLM-judges.
@@
  • We propose  [Uncaptioned image] MEDAL (Figure 1) for building multilingual meta-evaluation benchmarks, with modular generation, labelling, and curation stages [^3] While not the focus of this work, the large-scale raw dialogues and labels produced during MEDAL’s generation and automated annotation stages can be repurposed to pre-train lightweight evaluators or to fine-tune dialogue agents..
  • An automated analysis of chatbot responses shows that there are significant cross-lingual performance differences between different open-access LLMs.
  • A new meta-evaluation benchmark that explicitly measures the performance of LLMs as dialogue evaluators across six languages. Our analysis shows (1) a significant gap in performance between humans and LLMs, especially for more nuanced quality dimensions; and (2) reasoning models achieve better evaluative performance.
@@ 2.1 LLM-driven Dataset Curation
Several studies suggest leveraging LLM’s extensive world knowledge and linguistic knowledge for augmentation and synthetic data generation in order to scale dataset sizes (Ding et al., 2023; Li et al., 2023; Long et al., 2024). LLMs have been used to scale datasets for classification (Sahu et al., 2022; Hartvigsen et al., 2022), dialogue (Chen et al., 2023; Kim et al., 2023), and evaluation (Perez et al., 2022; Sorensen et al., 2024; Chiang and Lee, 2023; Tan et al., 2024).
@@ 3 Dialogue Generation
To mitigate the risks of synthetic data, we adopt a multi-agent setup where we recruit several LLMs acting as the user (§3.1.1) and employ a response generation feedback loop using a LLM as a user judge (§3.1.2) which is externally validated by human annotators (§3.4).
@@ 3.1.1 LLM as a User
To maximise diversity and minimise inherit biases, we condition the generation of dialogues using several contextual cues, including a scene description obtained from a commonsense knowledge graph, persona, an affective state and a target language. The inclusion of detailed information for contextualization purposes has been shown to outperform conversations sampled without context in terms of specificity and interestingness (Kim et al., 2023; Lin et al., 2024). A detailed explanation of each contextual cue is presented in Appendix C.
@@ 3.3 Experiments
We use GPT-4o-mini5552024-07-18 snapshot accessed March 2025. (OpenAI et al., 2024) and Gemma-3-27b-it (Gemma et al., 2025) as our users, and Gemini-2.0-Flash (Google, 2025b) as the online evaluator, given the strong multilingual capabilities and reasonable size/price of these models. For chatbot responses, we employ 8 open access LLMs of 4 different families: (1) Aya Expanse (Dang et al., 2024) with sizes 8B and 32B; (2) LLama-3.1 Instruct (Grattafiori et al., 2024) with sizes 8B and 70B; (3) Qwen2.5-Instruct (Qwen et al., 2025) with sizes 3B, 7B, and 72B; (4) and finally Phi-3.5-mini-Instruct (Abdin et al., 2024). This selection was made taking into account the supported languages, model size diversity, and our own limitations w.r.t to infrastructure.
@@
Automated Filtering.
Despite the inclusion of a feedback loop, an initial small-scale evaluation experiment uncovered instances where the user LLM did not adhere to the specified language or failed to maintain its role as a human user (i.e by erroneously reversing the role and acting as the responding chatbot). We therefore employ an additional filtering step using GPT-4.1 to automatically detect and remove these dialogues, while taking care to keep dialogues that exhibit chatbot deficiencies.
Statistics
In total, 38,400 dialogues were generated, stemming from 4,800 unique conversation starters (400 scenes, 2 user LLMs, and 6 languages). The automated filtering step then removed 6.44% of the dialogues, resulting in a final dataset of 35,927 instances. Additional statistics such as average turn length and diversity metrics are included in Appendix C.
3.4 Human Validation.
In order to evaluate the performance of the LLM as a user, we conduct a small scale human validation, where we ask annotators to rate the quality of the utterances generated by the LLM acting as a user in terms of their human-likeness, on a 1-5 Likert scale. Our annotator pool consists of a mix of graduate professionals in linguistics and computational linguistics, all with prior experience with annotations [^6] Finch et al. (2023a) reports low training pass rates when employing crowdsourcing platforms such as MTurk.. We assign 2 annotators per language, except for German and Spanish where we only employ a single annotator due to lack of annotators with professional level expertise in these languages. In total, 100 dialogues per language were evaluated. Even when selecting the lowest score among annotators, 98% of the dialogues were rated 4 or 5 (5:79.3%; 4:18.7%; 3:1.5%; 2:0.3%; 1:0%), with an average agreement (measured with Krippendorff’s $\alpha$) of .2592, denoting fair agreement.
4 Chatbot Evaluation
@@ 4.1 Experimental Setup
We evaluate the performance of the chatbots with GPT-4.1. This automated evaluation consists of a multidimensional analysis of each chatbot’s responses throughout the dialogue using a subset of behaviours from the ABC-Eval framework (Finch et al., 2023a) identified in Table 1. More details regarding this automated analysis, including the full prompt used for instructing the LLM is available in Appendix D.
@@ 4.2 Results
Figure 2 presents the percentage of detected issues across the 8 dimensions (uninterpretable, unsafe, lacks empathy, lacks commonsense, repetitive, incoherent, irrelevant, and non-factual), together with the average overall score for all tested chatbots. Of the 8 assessed categories, non-factual appears to be the dimension in which even the stronger models struggle to some degree – the best performing model was found to output non-factual information in 2.7% of its dialogues.
@@ 5.1 Benchmark Curation
In order to construct a meta-evaluation benchmark that maximizes the diversity and representativeness of dialogue quality, we employed a multi-stage, balanced sampling strategy to select 100 dialogues per target language. With a total of 600 dialogues averaging 9 turns per dialogue, our benchmark’s size is larger than other influential dialogue-level benchmarks [^7] DSTC11 (Rodríguez-Cantelar et al., 2023) evaluates 277 dialogues, FED (Mehri and Eskenazi, 2020a) 120 dialogues and ABC-Eval (Finch et al., 2023b) 400 dialogues..
@@ 5.2 Human Annotations
The annotators recruited to conduct the evaluation of the human-likeness of user generated utterances (Section 3) were also tasked with evaluating the performance of the chatbot. Each annotator was asked to identify whether a given issue was present in the chatbot responses (binary classification), and in the end provide an overall assessment of the chatbot (1-5) throughout the dialogue. Instructions emphasized the need to evaluate dialogues as if they were online chit-chat interactions.
Agreement values (reported using Krippendorff’s $\alpha$) are generally within the agreement reported by Finch et al. (2023a) for behaviour labels ($0.3\leq\alpha\leq 0.8$). Inter-annotator agreement was substantial for objective dimensions such as uninterpretable and non-factual, but lower for lacks empathy and repetitive which are more nuanced. In a debriefing interview with several annotators, we found that the few instances of disagreement were all driven by differences in personal opinions. This highlights the value of human annotation in capturing a diverse range of perspectives (Plank, 2022). Additional details regarding this annotation are presented in Appendix E.1.
@@ 5.3 Meta-Evaluation of LLMs as Automated Dialogue Evaluators
5.4 Meta-evaluation Results
@@
Looking at F1- performance (Figure 5), all evaluated models achieve results comparable to human expert performance. This indicates that identifying issue-free dialogues is a relatively easy task for current LLMs. The F1+ results, however, reveal a significant gap between human and model performance across most failure categories. While human annotators retain strong F1+ scores across all categories, the evaluated models struggle to match this performance.
@@ 5.4.1 Discussions
When comparing the performance of the reasoning models against the non-reasoning models we note that the reasoning models showcase better performance on both tasks, thanks for the most part due to better recall. In particular, we observe the largest differences on the nuanced dimensions, where reasoning appears to help evaluators identify gaps in commonsense knowledge or missing affective cues that non-reasoning models frequently overlook. The exception to this is GPT-4.1, where it generally matches o3 performance.
@@
Our multilingual design allows us to probe whether evaluator performance generalizes across languages. While sample size and confounding factors such as annotator variation and uneven dialogue quality make language-wise results difficult to generalise, we nevertheless observe consistent trends (Figures 4 and 13). For example, GPT-4.1 and O3 achieve relatively stable performance across all six languages, while DeepSeek-V3 significantly underperforms in Spanish and French, indicating this model may suffer from training data imbalances or limited cross-lingual generalization for these languages.
6 Findings and Contributions
This work introduces  [Uncaptioned image] MEDAL, a novel framework for generating and curating simulated multilingual open-domain human-chatbot conversations in support of evaluating current LLMs as both chatbots and evaluators. MEDAL operates in stages: first, it scales multilingual synthetic dialogue generation via a multi-LLM approach, with LLM agents acting as the user, chatbot, and judge. MEDAL then employs a strong LLM-judge for large-scale automated evaluation, yielding insights into the performance of small-to-medium size open-access LLMs as chatbots. These inform the automated curation of our meta-evaluation benchmark, which is used to evaluate several state-of-the-art LLMs acting as dialogue evaluators.
Our key contributions and findings are:
  1. 1.
    MEDAL provides a novel framework for on-demand generation, evaluation and curation of multilingual dialogues, and evaluation benchmarks. MEDAL can be further extended to include additional languages and/or LLMs (as users, chatbots, or judges).
  2. 2.
    We conduct a comprehensive evaluation of LLMs as multilingual chatbots finding that there are significant cross-lingual performance differences between the primary open-access LLMs.
  3. 3.
    We benchmark LLMs as evaluators showing that models struggle with nuanced conversational dimensions (e.g., empathy, commonsense), and that reasoning models consistently outperform non-reasoning counterparts, particularly on these harder cases.
@@ C.2 Generation Details
We set a temperature of 0.1 and limit the generation to 64 tokens to force smaller feedback. Due to the creativity imposed by the generation parameters, we set the regeneration limit of the first User turn to 10 attempts, which was sufficient to have all seed contexts accepted by the LLM-judge. For subsequent turns, we limit regeneration attempts to 5, which achieves a 90% pass-rate, and keeps costs below a factor of 4101010Incorporating the feedback loop makes the process 4 times more expensive in terms of LLM queries, assuming response generation and evaluation steps have equal cost.. Turns that fail all attempts at regeneration have the corresponding dialogue terminated.
@@
Table 2: Statistics of the generated dataset compared to other open-domain dialogue datasets. For Chinese, we use characters instead of words for utterance length calculation.
C.3 Statistics

### v4

# John Mendonça1,2, Alon Lavie3,4    Isabel Trancoso1,2

@@ Abstract
Evaluating the quality of open-domain chatbots has become increasingly reliant on LLMs acting as automatic judges. However, existing meta-evaluation benchmarks are static, outdated, and lacking in multilingual coverage, limiting their ability to fully capture subtle weaknesses in evaluation. We introduce  [Uncaptioned image] MEDAL, an automated multi-agent framework for curating more representative and diverse open-domain dialogue evaluation benchmarks [^1] Code and data alongside submission.. Our approach leverages several LLMs to generate user-chatbot multilingual dialogues, conditioned on varied seed contexts. Then, a state-of-the-art LLM (GPT-4.1) is used for a multidimensional analysis of the performance of the chatbots, uncovering noticeable cross-lingual performance differences. Guided by this large-scale evaluation, we curate a new meta-evaluation multilingual benchmark and human-annotate samples with nuanced quality judgments. This benchmark is then used to assess the ability of several reasoning and non-reasoning LLMs to act as evaluators of open-domain dialogues. Using MEDAL, we uncover that state-of-the-art judges fail to reliably detect nuanced issues such as lack of empathy, common sense, or relevance.
@@
John Mendonça1,2, Alon Lavie3,4and Isabel Trancoso1,2 1 INESC-ID, Lisbon 2 Instituto Superior Técnico, University of Lisbon 3 Carnegie Mellon University, Pittsburgh 4 Phrase, Pittsburgh {john.mendonca, isabel.trancoso}@inesc-id.pt, alavie@cs.cmu.edu
@@ 1 Introduction
Figure 1: Overview of  Refer to caption MEDAL, a 3-step multi-agent framework for curating native multilingual dialogue meta-evaluation benchmarks. ➊ Dialogue generation (§3): Generation of multi-turn, multilingual synthetic dialogues seeded from diverse contexts, between an LLM acting as a user (with utterances validated by an LLM-judge within a feedback loop) and the target chatbot; ➋ Automated labelling (§4): Multidimensional evaluation of all dialogues by a strong LLM (GPT-4.1); ➌ Benchmark Curation (§5): Curation of meta-evaluation benchmarks using human annotations used to assess LLM judges.
@@
    Dialogue generation (§3): synthetically generate a large collection of dialogues in multiple languages, using several LLMs acting as users and chatbots.
@@
    Benchmark curation (§5): guided by these automated labels, we curate and human-annotate a balanced set of dialogues to form a meta-evaluation benchmark that directly assesses LLM-judges.
@@
  • We propose  [Uncaptioned image] MEDAL (Figure 1) for building multilingual meta-evaluation benchmarks, with modular generation, labelling, and curation stages [^3] While not the focus of this work, the large-scale raw dialogues and labels produced during MEDAL’s generation and automated annotation stages can be repurposed to pre-train lightweight evaluators or to fine-tune dialogue agents.. MEDAL can be further extended to include additional languages and/or LLMs (as users, chatbots, or judges).
  • We conduct a comprehensive evaluation of LLMs as multilingual chatbots finding that there are significant cross-lingual performance differences between open-access LLMs.
  • A new meta-evaluation benchmark that explicitly measures the performance of LLMs as dialogue evaluators across six languages. Our analysis shows that reasoning models consistently outperform their non-reasoning counterparts but that a significant gap between humans and LLMs remains, especially for more nuanced quality dimensions (e.g., empathy, common sense).
@@ 2.1 LLM-driven Dataset Curation
Several studies suggest leveraging LLM’s extensive world knowledge and linguistic knowledge for augmentation and synthetic data generation in order to scale dataset sizes (Ding et al., 2023; Li et al., 2023; Long et al., 2024). LLMs have been used to scale datasets for classification (Sahu et al., 2022; Hartvigsen et al., 2022), dialogue (Chen et al., 2023; Kim et al., 2023), and evaluation (Perez et al., 2022; Chiang and Lee, 2023; Tan et al., 2024).
@@ 3 Dialogue Generation
To mitigate the risks of synthetic data, we adopt a multi-agent setup where we recruit several LLMs acting as the user (§3.1.1) and employ a feedback loop using an LLM as a user judge (§3.1.2). The resulting user turns are externally validated by human annotators (§3.5).
@@ 3.1.1 LLM as a User
To maximise diversity and minimise inherit biases, we condition the generation of dialogues using several contextual cues, including a scene description (obtained from a commonsense knowledge graph), a persona, an affective state and a target language. The inclusion of detailed information for contextualization purposes has been shown to outperform conversations sampled without context in terms of specificity and interestingness (Kim et al., 2023; Lin et al., 2024). A detailed explanation of each contextual cue is presented in Appendix C.
@@ 3.3 Experiments
We use GPT-4o-mini5552024-07-18 snapshot accessed March 2025. (OpenAI et al., 2024) and Gemma-3-27b-it (Gemma et al., 2025) as our users, and Gemini-2.0-Flash (Google, 2025b) as the user judge, given the strong multilingual capabilities and reasonable size/price of these models. For chatbot responses, we employ 8 open access LLMs of 4 different families: (1) Aya Expanse (Dang et al., 2024) with sizes 8B and 32B; (2) LLama-3.1 Instruct (Grattafiori et al., 2024) with sizes 8B and 70B; (3) Qwen2.5-Instruct (Qwen et al., 2025) with sizes 3B, 7B, and 72B; (4) and finally Phi-3.5-mini-Instruct (Abdin et al., 2024). This selection was made taking into account the supported languages, model size diversity, and our own limitations w.r.t to infrastructure.
@@
Statistics.
In total, 38,400 dialogues were generated, stemming from 4,800 unique conversation starters (400 scenes, two user LLMs, and six languages). Additional statistics such as average response length, number of turns and diversity metrics are included in Appendix C.
3.4 Qualitative Analysis
Despite the inclusion of a feedback loop, a cursory analysis uncovered instances where the user LLM failed to maintain its role as a human user (i.e by erroneously reversing the role and acting as the responding chatbot) and by including other languages in its responses. These issues highlight limitations of our LLM judge (Gemini-2.0-Flash) in reliably identifying conversational roles and maintaining consistent linguistic output across turns. Subsequently, we employ GPT-4.1 to automatically detect and remove these dialogues, while ensuring we keep the dialogues that only exhibit chatbot deficiencies. We find that 6.44% of dialogues were flagged as having malformed user responses (4.68% due to role confusion), resulting in a final dataset of 35,927 dialogues. Additional analysis are available in Appendix C.3.
3.5 Human Validation
In order to evaluate the performance of the LLM as a user, we conduct a small scale human validation, where we ask annotators to rate the quality of the utterances generated by the LLM acting as a user in terms of their human-likeness, on a 1-5 Likert scale. Our annotator pool consists of a mix of graduate professionals in linguistics and computational linguistics, all with prior experience with annotations [^6] Finch et al. (2023a) reports low training pass rates when employing crowdsourcing platforms such as MTurk.. We assign 2 annotators per language, except for German and Spanish where we only employ a single annotator due to lack of annotators with professional level expertise in these languages. We validated the quality of user responses in the 600 dialogues (100 per language) later used for multidimensional human annotation (see Section 5). Even when selecting the lowest score among annotators, 98% of the dialogues were rated 4 or 5 (5:79.3%; 4:18.7%; 3:1.5%; 2:0.3%; 1:0%), with an average agreement (measured with Krippendorff’s $\alpha$) of .2592, denoting fair agreement.
4 Automated Labelling
@@ 4.1 Experimental Setup
We evaluate the performance of the chatbots with GPT-4.1. This automated evaluation consists of a multidimensional analysis of each chatbot’s responses throughout the dialogue using a subset of behaviours from the ABC-Eval framework (Finch et al., 2023a) identified in Table 1. More details regarding this automated analysis, including the full prompt used for instructing the LLM are available in Appendix D.
@@ 4.2 Results
Figure 2 presents the percentage of detected issues across the 8 dimensions, together with the average overall score for all tested chatbots. Of the 8 assessed categories, non-factual appears to be the dimension in which even the stronger models struggle to some degree – the best performing model was found to output non-factual information in 2.7% of its dialogues.
@@ 5.1 Benchmark Curation
In order to construct a meta-evaluation benchmark that maximizes the diversity and representativeness of dialogue quality, we employed a multi-stage, balanced sampling strategy to select 100 dialogues per target language. With a total of 9,000 assessments (9 quality dimensions $\times$ 10 annotators $\times$ 100 dialogues) spanning 600 unique dialogues, our benchmark’s size is larger than other influential dialogue-level benchmarks [^7] DSTC11 (Rodríguez-Cantelar et al., 2023) evaluates 277 dialogues, FED (Mehri and Eskenazi, 2020a) 120 dialogues and ABC-Eval (Finch et al., 2023b) 400 dialogues..
@@ 5.2 Human Annotations
The annotators recruited to conduct the evaluation of the human-likeness of user generated utterances (Section 3) were also tasked with evaluating the performance of the chatbot on the same set of dialogues. Each annotator was asked to identify whether a given issue was present in the chatbot responses (binary classification), and in the end provide an overall assessment of the chatbot (1-5) throughout the dialogue. Instructions emphasized the need to evaluate the chatbots as if they were employed in online chit-chat interactions.
Inter-annotator agreement was high overall, with annotators agreeing on over 89% of cases on average across languages and dimensions (Table 3). Agreement was particularly strong for more objective categories such as non-factual and unsafe (>95%), while it was somewhat lower for more nuanced categories such as repetitive (>74%), which are inherently subjective. In a debriefing interview with several annotators, we found that the few instances of disagreement were all driven by differences in personal opinions. This highlights the value of human annotation in capturing a diverse range of perspectives (Plank, 2022). Additional details regarding this annotation are presented in Appendix E.1.
@@ 5.3 Meta-Evaluation of LLMs as Automated Dialogue Evaluators
5.3.1 Meta-evaluation Results
@@
In terms of F1- (Figure 5), all evaluated models achieve results comparable to human expert performance. This indicates that identifying issue-free dialogues is a relatively easy task for current LLMs. The F1+ results, however, reveal a significant gap between human and model performance across most failure categories. While human annotators retain strong F1+ scores across all categories, the evaluated models struggle to match this performance.
@@ 5.3.2 Discussions
When comparing the performance of the reasoning models against the non-reasoning models we note that the reasoning models showcase better performance on both the overall prediction and issue detection, thanks for the most part due to better recall. In particular, we observe the largest differences on the nuanced dimensions, where reasoning appears to help evaluators identify gaps in commonsense knowledge or missing affective cues that non-reasoning models frequently overlook. The exception to this is GPT-4.1, where it generally matches o3 performance.
@@
Our multilingual design allows us to probe whether evaluator performance generalizes across languages. While sample size and confounding factors such as annotator variation and uneven dialogue quality make language-wise results difficult to generalise, we nevertheless observe consistent trends (Figures 4 and 14). For example, GPT-4.1 and o3 achieve relatively stable performance across all six languages, while DeepSeek-V3 significantly underperforms in Spanish and French, indicating this model may suffer from training data imbalances or limited cross-lingual generalization for these languages.
6 Conclusions
This work introduces  [Uncaptioned image] MEDAL, a novel and scalable framework for generating and curating native multilingual open-domain dialogue meta-evaluation benchmarks. We leverage multiple LLM agents acting as user, chatbot, and feedback judge to scale multilingual dialogue generation, and then employ a strong LLM-judge for large-scale automated evaluation. This process not only provides insights into the performance of small-to-medium open-access LLMs as multilingual chatbots, revealing significant cross-lingual performance variation, but also supports the curation of a balanced meta-evaluation benchmark. Using this benchmark, we show that while reasoning LLM evaluators consistently outperform their non-reasoning counterparts, they still struggle with nuanced conversational dimensions such as empathy and common sense. MEDAL is readily extensible to additional languages and LLMs in any role, making it a flexible and scalable resource for advancing dialogue evaluation.
@@ C.2 Generation Details
We set a temperature of 0.1 and limit the generation to 64 tokens to force smaller feedback. Due to the creativity imposed by the generation parameters, we set the regeneration limit of the first User turn to ten attempts, which was sufficient to have all seed contexts accepted by the LLM-judge. For subsequent turns, we limit regeneration attempts to five, which achieves a 90% pass-rate, and keeps costs below a factor of four [^10] Incorporating the feedback loop makes the process 4 times more expensive in terms of LLM queries, assuming response generation and evaluation steps have equal cost.. Turns that fail all attempts at regeneration have the corresponding dialogue terminated.
@@
Table 2: Statistics of the generated dataset compared to other open-domain dialogue datasets. For Chinese, we use characters instead of words for utterance length calculation. We exclude the average calculation due to this reason.
C.3 Detection of Malformed User Behaviour
Figure 7: Proportion of dialogues automatically filtered due to role reversal or unintended language switching. Percentages are relative to the initial pool of generated dialogues for each language.
We used GPT-4.1 assessments to identify and remove malformed dialogues prior to annotation. These included two systematic error types observed during generation: (i) role confusion, where the user agent incorrectly assumed the role of the chatbot, and (ii) language mixing, where the chatbot produced responses in a language other than the intended target.
Across all six target languages and two user LLMs, 6.44% of dialogues were flagged as malformed (4.68% due to role confusion) and excluded from subsequent stages. Figure 7 reports the percentage of removed dialogues. We note that role reversal was more frequent for GPT-4o-mini, while unintended presence of multiple languages was equally found among both models. Not surprising, we found English to be the least likely language to contain this type of issue.
C.4 Statistics